A survey of 25 randomly selected customers found the ages shown (in years). The mean is 33.60 years and the standard deviation is 9.52 years a) Construct a 90% confidence interval for the mean age of all customers, assuming that the assumptions and conditions for the confidence interval have been met. b) How large is the margin of error? c) How would the confidence interval change if you had assumed that the standard deviation was known to be 10.0 years? 44 46 23 31 46 39 35 42 28 39 28 35 35 42 49 36 38 38 11 25 20 25 29 37 19
Answer #1

a) Margin of error = (\frac{1.7109*9.52}{\sqrt{25}})

= 3.26

90% Interval

(33.60 -3.26 , 33.60+3.26)

= (30.34, 36.86)

b) Margin of error is = 3.26

c) Margin of error = (\frac{1.7109*10}{\sqrt{25}})

= 3.42

90% Interval

(33.60-3.42, 33.60+3.42)

= (30.18, 37.02)